OpenStudy (anonymous):

Trig Identities Proof: \[\sin^2x+\cos^4x=\cos^2x+\sin^4x\]

OpenStudy (michele_laino):

hint: left side= \[\Large {\left( {\sin x} \right)^2} + {\left( {\cos x} \right)^4} = {\left( {\sin x} \right)^2} + {\left( {1 - {{\left( {\sin x} \right)}^2}} \right)^2} = ...?\] please continue

rvc (rvc):

yep now it's easy :) just expand it :)

OpenStudy (anonymous):

Okay! I got it! :) Thanks you guys! @Michele_Laino and @rvc

OpenStudy (michele_laino):

Thank you! :) @esam2

rvc (rvc):

thank you